OpenStudy (nehemias123):

What is the value of r in the equation r + 9.2 = 17.5?

OpenStudy (nehemias123):

need hlep

OpenStudy (firejay5):

17.5 - 9.2 = 8.3, so therefore, r = 8.3

OpenStudy (firejay5):

to correctly solve it, subtract 9.5 from both sides of the equation to get r by itself

OpenStudy (nehemias123):

so it is 8.3

OpenStudy (anonymous):

In order to get the value of r, you have to isolate the r by subtracting 9.2 from 17.5. You subtract 9.2 from both sides to isolate the r. leaving r= 17.5 - 9.2

OpenStudy (firejay5):

exactly the correct answer is 8.3

OpenStudy (anonymous):

Does 17.5 - 9.2 equal 8.3? If it does, than yes @NEHEMIAS123

OpenStudy (xapproachesinfinity):

hmm let's see, i would say 8.3 as well

OpenStudy (nehemias123):

ok thak you all

sammixboo (sammixboo):

Make sure not to give out the answers! But this is pretty easy to solve =) Got to subtract 9.2 on both sides of the equation r + 9.2 - 9.2 = 17.5 - 9.2 r = 17.5 - 9.2